# Imprints of the Local Bubble and Dust Complexity on Polarized Dust Emission

### v1

@@ Abstract
Using 3D dust maps and Planck polarized dust emission data, we investigate the influence of the 3D geometry of the nearby interstellar medium (ISM) on the statistics of the dust polarization on large ($80^{\prime}$) scales. We test the idea that the magnetic field in the nearby dust is preferentially tangential to the Local Bubble wall, but we do not find an imprint of the Local Bubble geometry on the dust polarization fraction. We also test the hypothesis that the complexity of the 3D dust distribution drives some of the measured variation of the dust polarization fraction. We compare sightlines with similar total column densities and find evidence that, on average, the dust polarization fraction decreases when the dust column is substantially distributed among multiple components at different distances. Conversely, the polarization fraction is higher for sightlines where the dust is more concentrated in 3D space. This finding is statistically significant for the dust within 1.25 kpc, but the effect disappears if we only consider dust within 270 pc. The extended 3D dust distribution, rather than solely the dust associated with the Local Bubble, plays a role in determining the observed dust polarization fraction on these scales. This conclusion is consistent with a simple analytical prediction and remains robust under various modifications to the analysis. These results illuminate the relationship between the 3D geometry of the ISM and tracers of the interstellar magnetic field. We also discuss implications for our understanding of the polarized dust foreground to the cosmic microwave background.
@@ 1 Introduction
Interstellar magnetic fields play an important role in various astrophysical processes (see, e.g., Ferrière, 2001; Heiles & Haverkorn, 2012; Pattle et al., 2023). However, little is known about the magnetic field structure in the nearby interstellar medium (ISM). Some work has suggested connections between the local magnetic field structure and other tracers of ISM morphology, perhaps due to dynamical influences, e.g. the formation of structures like superbubbles (e.g., Santos et al., 2011; Frisch et al., 2012; Berdyugin et al., 2014; Tahani et al., 2022a, b).
@@
Recent evidence suggests that variations in the dust polarization fraction over large angular scales in the diffuse sky are mainly driven by the structure of the magnetic field. Planck Collaboration et al. (2020a) probed the influence of the magnetic field geometry on the dust polarization fraction by comparing the local polarization angle dispersion with the 353 GHz polarization fraction. They calculated the polarization angle dispersion for an annulus with inner and outer radii of 40 and 120, and found that the 353 GHz polarization fraction is anti-correlated with the local polarization angle dispersion at 160 resolution. They showed that this relationship is consistent with models that only include topological effects of the turbulent magnetic field, but otherwise have uniform dust properties and alignment. Planck Collaboration et al. (2020a) conclude that the polarization fraction and the dispersion of polarization angles are similarly sensitive to the structure of the magnetic field. Hensley et al. (2019) further showed that some of the variability in the dispersion of polarization angles, and thus the polarization fraction, can be explained by the magnetic inclination angle, i.e., the angle between the magnetic field and the plane of the sky. The polarization fraction is maximized when the magnetic field is tangential to the plane of the sky and zero when it is parallel to the line of sight. Chen et al. (2019) and Sullivan et al. (2021) used statistical properties of the observed polarization fraction to estimate the average inclination angle of molecular cloud-scale magnetic fields.
@@
This paper focuses on latitudes and angular scales where we expect that the dust polarization primarily traces the nearby dust. Our Sun’s current location is near the center of a superbubble, which is thought to have been created by supernova explosions within the past 10-15×${}\times~{}10^{6}$ years (Cox & Reynolds, 1987; Maíz-Apellániz, 2001; Breitschwerdt et al., 2016). It is commonly known as the Local Bubble, Local Cavity, or Local Chimney (Welsh et al., 2004; Puspitarini & Lallement, 2012). Since star formation tends to be concentrated, sequential supernovae are common (Zucker et al., 2022; Watkins et al., 2023; Barnes et al., 2023; Sandstrom et al., 2023). Supernova explosions sweep up matter and magnetic field lines, leaving behind low-density superbubbles on the order of hundreds of parsecs in diameter (Kim & Ostriker, 2015). The swept-up matter is compressed into a shell surrounding the expanding superbubble, which is thought to trigger the formation of dense gas and stars (Elmegreen, 2011; Dawson, 2013; Inutsuka et al., 2015).
It is reasonable to expect that the formation of the Local Bubble dramatically influenced the magnetic field geometry in the nearby ISM. Some studies have aimed at modeling the geometry of the wall of cold neutral gas and dust surrounding the Local Bubble (e.g., Alves et al., 2018; Pelgrims et al., 2020). Since the geometry of the magnetic field affects the measured polarization fraction, we search for an imprint of the Local Bubble geometry on the dust polarization fraction in this paper.
Additionally, one of the probes of the 3D spatial distribution of the neutral ISM is dust extinction toward stars. This is due to the scattering and absorption of starlight by dust. The extinction of a star’s apparent magnitude is correlated with the dust column density along the line of sight from the observer to that star. The Gaia survey provided accurate distances to more than a billion stars within a few kiloparsecs from the Sun. Combining this distance information with the level of extinction towards each star has been transformative for the construction of 3D maps of the differential dust extinction (Lallement et al., 2019; Leike et al., 2020; Vergely et al., 2022; Edenhofer et al., 2023). We use several 3D dust maps to quantify the complexity of the spatial distribution of the dust along the line of sight. We use that to explore the relationship between the 3D dust distribution and the measured polarization fraction.
In this work, we investigate the relationship between the 3D geometry of the nearby ISM and the dust polarization fraction. We start by introducing the data we use in Section 2. In Section 3, we search for an imprint of the Local Bubble geometry on the dust polarization fraction. In Section 4, we test how the polarization fraction is affected by the line-of-sight complexity of the dust. We discuss the implications of our results and conclude in Section 5.
@@ 2 Data
Figure 2: Slices through the XY (left), XZ (middle), and YZ (right) planes of the 3D reconstructed differential extinction maps of Lallement et al. (2019) (top) and Edenhofer et al. (2023) (bottom). The Sun is at the origin. The positive X axis points towards the Galactic center at $l~{}=~{}0^{\circ}$, the positive Y axis points towards $l~{}=~{}90^{\circ}$ in the Galactic plane, and the positive Z axis points out of the plane in the direction of the Galactic North pole. The slices show the log scale of the differential extinction, which is in units of magnitudes per parsec. The subpanels only extend up to 400 pc in each direction for a direct comparison. The model for the Local Bubble surface geometry of Pelgrims et al. (2020) is overplotted in white in each subpanel.
@@ 2.1 Planck Data Products
We use the modified asymptotic estimator of Plaszczynski et al. (2014) to debias the polarized intensity of the GNILC dust emission map and the associated uncertainty. We obtain a signal-to-noise ratio map of the polarization fraction, SNRp, shown in Figure 1.
@@
In Section 3, we use the 3D model of the Local Bubble surface geometry constructed by Pelgrims et al. (2020). To create this model, Pelgrims et al. (2020) fits a shell of higher dust density around the Sun using the 3D Cartesian map of dust differential extinction constructed by Lallement et al. (2019). The Lallement et al. (2019) dust map is based on data from Gaia DR2 (Gaia Collaboration et al., 2018) and 2MASS (Skrutskie et al., 2006) and spans $6~{}\times~{}6~{}\times~{}0.8~{}{\rm kpc}^{3}$ in the Heliocentric right-handed Galactic-XYZ coordinates. It has a voxel volume of 125 pc3 and a spatial resolution of 25 pc. Example slices of this map with the Pelgrims et al. (2020) model overplotted are shown in Figure 2.
We also use the 3D model of the Local Bubble surface geometry constructed by O’Neill et al. (2024) for a brief investigation in Section 3. They use the 3D dust differential extinction provided by Edenhofer et al. (2023) to fit the distance to the Local Bubble in all directions as a region of higher dust density around the Sun.
In Section 4, we query the 12 posterior samples for the 3D dust maps provided by Edenhofer et al. (2023) via the publicly available dustmaps Python package (Green, 2018) at their plane-of-sky native angular resolution of $14^{\prime}$, which corresponds to a HEALPix pixelization scheme at $N_{\rm side}~{}=~{}{256}$. These 3D dust maps leverage distance and extinction estimates to stars from Zhang et al. (2023), which are derived from Gaia DR3 data (Gaia Collaboration et al., 2023). The distance resolution of these maps varies from 0.4 pc at 69 pc to 7 pc at 1.25 kpc. We query the map using uniform distance bins of 7 pc. The map is provided in unitless extinction values defined in Zhang et al. (2023). We multiply the map by a factor of 2.8 to convert it to Johnson’s $V$-band $A^{\prime}_{V}$ (Zhang et al., 2023). We then convert $A^{\prime}_{V}$ to volume density of hydrogen nuclei ($n_{\rm H}$) using the extinction curve from Fitzpatrick et al. (2019) to convert $A_{G}~{}=~{}0.796~{}A_{V}$ and the relationship $A_{G}~{}/~{}N_{\rm H}~{}=~{}4~{}\times~{}10^{-22}~{}{\rm cm}^{2}~{}{\rm mag}$ from Zucker et al. (2021) and Bialy et al. (2021). To match the resolution of the polarization fraction and smooth out small-scale fluctuations in the map, we smooth the HEALPix sphere at each distance bin to a FWHM= $80^{\prime}$ then repixelate it to $N_{\rm side}~{}=~{}{64}$. Example slices of the map of the mean of the posterior samples are also shown in Figure 2 with the Pelgrims et al. (2020) model overplotted for comparison.
@@
2.3 Rotation Measure
We use a map of the posterior mean of the Galactic Faraday sky developed by Hutschenreuter et al. (2023) for a brief investigation in Section 3. To produce these maps, they use information field theory, a Bayesian inference framework for fields (Enßlin, 2019), on compiled catalogs of extragalactic Faraday rotation measures (Van Eck et al., 2023) and Galactic pulsar dispersion measures (Manchester et al., 2005), as well as data on Galactic bremsstrahlung emission (Planck Collaboration et al., 2016b) and the hydrogen $\alpha$ spectral line (Finkbeiner, 2003).
@@ 3.1 Motivation
We begin the exploration of the effect of different geometrical factors on the polarization fraction by searching for an imprint of the geometry of the dust wall surrounding the Local Bubble on the Planck 353 GHz dust polarization fraction. In this subsection, we discuss the assumptions made in previous studies regarding the Local Bubble surface. These assumptions help us design a test for studying this effect in the next subsection. A significant detection of an imprint of the Local Bubble geometry on the polarization fraction would validate these assumptions.
@@
The first assumption is that the observed polarized dust emission is dominated by dust in the Local Bubble wall at the relevant angular scales and Galactic latitudes. This assumption, made in several analyses (e.g., Alves et al., 2018; O’Neill et al., 2023), is supported by several studies using optical starlight polarization data (Leroy, 1999; Andersson & Potter, 2006; Santos et al., 2011; Frisch et al., 2015; Medan & Andersson, 2019; Cotton et al., 2019). The alignment of neutral hydrogen structures at local velocities with starlight polarization toward stars at distances within a few hundred parsecs is consistent with a picture where most of these structures are positioned at comparable distances within the Local Bubble at high Galactic latitudes ($|b|~{}>~{}30^{\circ}$) (Clark et al., 2014). Gontcharov & Mosenkov (2019) observed that starlight polarization fraction plateaus after 150-250 pc across the sky. By comparing the 353 GHz polarized emission with the polarized optical starlight, Skalidis & Pelgrims (2019) find that most of the 353 GHz polarized emission signal is captured within the first 250 pc at $|b|~{}>~{}60^{\circ}$, suggesting the presence of a dust wall around that distance.
@@
The second assumption is that the magnetic field’s inclination is tangential to the surface of the Local Bubble. To fit a model of the Local Bubble magnetic field, Alves et al. (2018) assume that all the swept-up matter and field lines due to the supernova explosions that formed the Local Bubble are squeezed into a thin layer that follows its surface, leading the magnetic field lines to be tangent to the surface. O’Neill et al. (2023) also make this assumption to project the observed dust polarization angles onto the Bubble’s surface and build a 3D model of the Bubble wall magnetic field.
@@
Other work has found magnetic field structure tangential to bubbles on supernova scales, observationally (e.g., Kothes & Brown, 2009; Tahani et al., 2022a, b) and in simulations (Kim & Ostriker, 2015). It has also been shown on the scales of Hii regions, observationally (Tahani et al., 2023) and in simulations (Krumholz et al., 2007). On the scale of superbubbles, such as our Local Bubble, a comparison of the plane-of-the-sky and line-of-sight magnetic field strengths as well as measurements of the polarization fraction towards regions associated with the Orion-Eridanus superbubble suggest that the large-scale magnetic field in the region was primarily shaped by the expanding superbubble and is tangential to its surface (Heiles, 1997; Soler et al., 2018). However, there is no direct evidence that the nearby magnetic field is preferentially tangential to the Local Bubble surface. The only constraints on the magnetic inclination angle using the geometry of the Local Bubble surface are at the Galactic polar caps from Alves et al. (2018) and Pelgrims et al. (2020). They each fit an analytical parametric model of the magnetic field, including terms related to the geometry of the Bubble wall, to dust polarization data from Planck. One of the parameters of the magnetic field model used by Pelgrims et al. (2020) is the magnetic inclination angle. They find that the magnetic field is oriented mostly in the plane of the sky at the Galactic polar caps. The polar caps are the regions in which the Local Bubble has perhaps the least-well-defined surface, as discussed further below.
@@
The shape of the Local Bubble wall has been modeled differently in different works. Some studies fit a generalized parametric geometry, such as an ellipsoid (Alves et al., 2018), in an attempt to fit general properties of the Local Bubble magnetic field on large scales. Others model the detailed boundary of the Local Bubble using 3D maps of the dust extinction (Pelgrims et al., 2020; O’Neill et al., 2024). These models therefore vary based on the variations of the different 3D dust reconstructions. Pelgrims et al. (2020) model the radial distance of the Local Bubble wall from the Sun in each direction as the first distance where the second derivative with respect to the distance of the differential extinction constructed by Lallement et al. (2019) reaches zero, i.e., the first inflection point, ${\rm d}^{2}A_{V}^{\prime}(r)/{\rm d}r^{2}~{}=~{}0$. O’Neill et al. (2023) and Zucker et al. (2022) use the geometry defined by Pelgrims et al. (2020) in their work. O’Neill et al. (2024) employ a similar methodology to Pelgrims et al. (2020), using the differential extinction maps constructed by Edenhofer et al. (2023) instead to construct their model. Other studies use different tracers to model the geometry of the Bubble. Liu et al. (2017) assumes that the measured X-ray intensity is proportional to the distance to the Bubble in the considered direction. Several other tracers such as NaI absorption measurements (Sfeir et al., 1999; Lallement et al., 2003), stellar color excess measurements (Lallement et al., 2014), and diffuse interstellar bands (Farhang et al., 2019), have also been used for constructing models of the Bubble wall geometry. These geometries vary significantly from one another. While some model the Local Bubble as a closed surface (Pelgrims et al., 2020), others describe the same structure as a Local Chimney, i.e., open at high Galactic latitudes and funneling material into the Milky Way’s halo (e.g., Sfeir et al., 1999; Lallement et al., 2003; O’Neill et al., 2024). Also, the Local Bubble surface may have tunnels to surrounding cavities like the Gum Nebula and/or GSH238+00+09 rather than having a closed geometry (e.g., Welsh, 1991; Lallement et al., 2003; O’Neill et al., 2024).

### v2

@@ Abstract
Using 3D dust maps and Planck polarized dust emission data, we investigate the influence of the 3D geometry of the nearby interstellar medium (ISM) on the statistics of the dust polarization on large ($80^{\prime}$) scales. We test recent models that assume that the magnetic field probed by the polarized dust emission is preferentially tangential to the Local Bubble wall, but we do not find an imprint of the Local Bubble geometry on the dust polarization fraction. We also test the hypothesis that the complexity of the 3D dust distribution drives some of the measured variation of the dust polarization fraction. We compare sightlines with similar total column densities and find that, on average, the dust polarization fraction decreases when the dust column is substantially distributed among multiple components at different distances. Conversely, the dust polarization fraction is higher for sightlines where the dust is more concentrated in 3D space. This finding is statistically significant for the dust within 1.25 kpc, but the effect disappears if we only consider dust within 270 pc. In conclusion, we find that the extended 3D dust distribution, rather than solely the dust associated with the Local Bubble, plays a role in determining the observed dust polarization fraction at 80. This conclusion is consistent with a simple analytical prediction and remains robust under various modifications to the analysis. These results illuminate the relationship between the 3D geometry of the ISM and tracers of the interstellar magnetic field. We discuss implications for our understanding of the polarized dust foreground to the cosmic microwave background.
@@ 1 Introduction
Interstellar magnetic fields play an important role in various astrophysical processes (see, e.g., Ferrière, 2001; Heiles & Haverkorn, 2012; Pattle et al., 2023). However, little is known about the magnetic field structure in the nearby interstellar medium (ISM). Some works have suggested connections between the local magnetic field structure and other tracers of ISM morphology, perhaps due to dynamical influences, e.g. the formation of structures like superbubbles (e.g., Santos et al., 2011; Frisch et al., 2012; Berdyugin et al., 2014; Tahani et al., 2022a, b).
@@
Recent evidence suggests that variations in the fractional polarization of the dust emission, i.e., the ratio of the polarized to the total intensity of the dust emission, over large angular scales in the diffuse sky are mainly driven by the structure of the magnetic field. Henceforth, we will refer to the fractional polarization of the dust emission as the dust polarization fraction. Planck Collaboration et al. (2020a) probed the influence of the magnetic field geometry on the dust polarization fraction by comparing the local polarization angle dispersion with the 353 GHz polarization fraction. They calculated the polarization angle dispersion for an annulus with inner and outer radii of 40 and 120, and found that the 353 GHz polarization fraction is anti-correlated with the local polarization angle dispersion at 160 resolution. They showed that this relationship is consistent with models that only include topological effects of the turbulent magnetic field, but otherwise have uniform dust properties and alignment. Planck Collaboration et al. (2020a) conclude that the dust polarization fraction and the dispersion of polarization angles are similarly sensitive to the structure of the magnetic field. Hensley et al. (2019) further showed that some of the variability in the dispersion of polarization angles, and thus the dust polarization fraction, can be explained by the magnetic inclination angle, i.e., the angle between the magnetic field and the plane of the sky. The dust polarization fraction is maximized when the magnetic field is tangential to the plane of the sky and zero when it is parallel to the line of sight. Chen et al. (2019) and Sullivan et al. (2021) used statistical properties of the observed dust polarization fraction to estimate the average inclination angle of molecular cloud-scale magnetic fields.
@@
Our Sun’s current location is near the center of a superbubble, which is thought to have been created by supernova explosions within the past 10-15×${}\times~{}10^{6}$ years (Cox & Reynolds, 1987; Maíz-Apellániz, 2001; Breitschwerdt et al., 2016; Schulreich et al., 2023). It is commonly known as the Local Bubble, Local Cavity, or Local Chimney (Welsh et al., 2004; Puspitarini & Lallement, 2012). Since star formation tends to be concentrated, sequential supernovae are common (Zucker et al., 2022; Watkins et al., 2023; Barnes et al., 2023; Sandstrom et al., 2023). Supernova explosions sweep up matter and magnetic field lines, leaving behind low-density superbubbles on the order of hundreds of parsecs in diameter (Kim & Ostriker, 2015). The swept-up matter is compressed into a shell surrounding the expanding superbubble, which is thought to trigger the formation of dense gas and stars (Elmegreen, 2011; Dawson, 2013; Inutsuka et al., 2015).
Thus, it is reasonable to expect that the formation of the Local Bubble dramatically influenced the magnetic field geometry in the nearby ISM. Some studies have aimed at modeling the geometry of the wall of cold neutral gas and dust surrounding the Local Bubble (e.g., Alves et al., 2018; Pelgrims et al., 2020). Since the geometry of the magnetic field affects the measured dust polarization fraction, we search for an imprint of the Local Bubble geometry on the dust polarization fraction in this paper.
Additionally, one of the probes of the 3D spatial distribution of the neutral ISM is dust extinction toward stars. This is due to the scattering and absorption of starlight by dust. The extinction of a star’s apparent magnitude is correlated with the dust column density along the line of sight from the observer to that star. The Gaia survey provided accurate distances to more than a billion stars within a few kiloparsecs from the Sun. Combining this distance information with the level of extinction towards each star has been transformative for the construction of 3D maps of the differential dust extinction (Lallement et al., 2019; Leike et al., 2020; Vergely et al., 2022; Edenhofer et al., 2023). We use several 3D dust maps to quantify the complexity of the spatial distribution of the dust along the line of sight. We use that to explore the relationship between the 3D dust distribution and the measured dust polarization fraction.
In this work, we investigate the relationship between the 3D geometry of the nearby ISM and the dust polarization fraction. We start by introducing the data we use in Section 2. In Section 3, we search for an imprint of the Local Bubble geometry on the dust polarization fraction. In Section 4, we test how the dust polarization fraction is affected by the line-of-sight complexity of the dust. We discuss the implications of our results and conclude in Section 5.
@@ 2 Data
Figure 2: Slices through the XY (left), XZ (middle), and YZ (right) planes of the 3D reconstructed differential extinction maps of Lallement et al. (2019) (top) and Edenhofer et al. (2023) (bottom). The Sun is at the origin. The positive X axis points towards the Galactic center at $l~{}=~{}0^{\circ}$, the positive Y axis points towards $l~{}=~{}90^{\circ}$ in the Galactic plane, and the positive Z axis points out of the plane in the direction of the Galactic North pole. The slices show the log of the differential extinction, which is in units of magnitudes per parsec. The subpanels only extend up to 400 pc in each direction for a direct comparison. The model for the Local Bubble surface geometry of Pelgrims et al. (2020) is overplotted in white in each subpanel. The white disk at the center of the bottom panel is due to missing data within 70 pc of the Sun in the Edenhofer et al. (2023) maps.
@@ 2.1 Planck Data Products
We use the modified asymptotic estimator of Plaszczynski et al. (2014) to debias the polarized intensity of the GNILC dust emission map and the associated uncertainty. We obtain a signal-to-noise ratio map of the dust polarization fraction, SNRp, shown in Figure 1.
@@
In Section 3, we use the 3D model of the Local Bubble surface geometry constructed by Pelgrims et al. (2020). To create this model, Pelgrims et al. (2020) extract distances to the first high dust density regions around the Sun from the 3D Cartesian map of dust differential extinction constructed by Lallement et al. (2019). They smooth the map of the distances to the Local Bubble surface by filtering out spherical harmonic modes above some threshold to remove small-scale fluctuations that might appear due to an inhomogeneous distribution of the dust density on the small scales. In this work, we use the map filtered to $\ell_{\rm max}=10$, which is the map used in their analysis. The Lallement et al. (2019) dust map is based on data from Gaia DR2 (Gaia Collaboration et al., 2018) and 2MASS (Skrutskie et al., 2006) and spans $6~{}\times~{}6~{}\times~{}0.8~{}{\rm kpc}^{3}$ in the Heliocentric right-handed Galactic-XYZ coordinates. It has a voxel volume of 125 pc3 and a spatial resolution of 25 pc. Example slices of this map with the Pelgrims et al. (2020) model overplotted are shown in Figure 2.
We also use the 3D model of the Local Bubble surface constructed by O’Neill et al. (2024) for a brief investigation in Section 3. They use the 3D dust differential extinction provided by Edenhofer et al. (2023) to extract the distance to the Local Bubble in all directions as a region of higher dust density around the Sun.
In Section 4, we query the 12 posterior samples for the 3D dust maps provided by Edenhofer et al. (2023) via the publicly available dustmaps Python package (Green, 2018) at their plane-of-sky native angular resolution of $14^{\prime}$, which corresponds to a HEALPix pixelization scheme at $N_{\rm side}~{}=~{}{256}$. These 3D dust maps leverage distance and extinction estimates to stars from Zhang et al. (2023), which are derived from Gaia DR3 data (Gaia Collaboration et al., 2023). The distance resolution of these maps varies from 0.4 pc at 69 pc to 7 pc at 1.25 kpc. We query the map using uniform distance bins of 7 pc. The map is provided in unitless extinction values defined in Zhang et al. (2023). We multiply the map by a factor of 2.8 to convert it to Johnson’s $V$-band $A^{\prime}_{V}$ (Zhang et al., 2023). We then convert $A^{\prime}_{V}$ to volume density of hydrogen nuclei ($n_{\rm H}$) using the extinction curve from Fitzpatrick et al. (2019) to convert $A_{G}~{}=~{}0.796~{}A_{V}$ and the relationship $A_{G}~{}/~{}N_{\rm H}~{}=~{}4~{}\times~{}10^{-22}~{}{\rm cm}^{2}~{}{\rm mag}$ from Zucker et al. (2021) and Bialy et al. (2021). To match the resolution of the dust polarization fraction and smooth out small-scale fluctuations in the map, we smooth the HEALPix sphere at each distance bin to a FWHM= $80^{\prime}$ then repixelate it to $N_{\rm side}~{}=~{}{64}$. Example slices of the raw map of the mean of the posterior samples are also shown in Figure 2 with the Pelgrims et al. (2020) model overplotted for comparison.
@@
2.3 Galactic Faraday Rotation Measure
We use the all-sky Galactic Faraday rotation measure (RM) map produced by Hutschenreuter et al. (2023) for a brief investigation in Section 3. Using information field theory, a Bayesian inference framework for fields (Enßlin, 2019), Hutschenreuter et al. (2023) disentangle the Galactic contribution to the RM from the compiled RM catalogs of polarized radio sources such as radio galaxies (Van Eck et al., 2023), supplemented by Galactic pulsar dispersion measures (Manchester et al., 2005), as well as data on Galactic bremsstrahlung emission (Planck Collaboration et al., 2016b) and the hydrogen $\alpha$ spectral line (Finkbeiner, 2003).
@@ 3.1 Motivation
We begin the exploration of the effect of different geometrical factors on the dust polarization fraction by searching for an imprint of the geometry of the dust wall surrounding the Local Bubble on the Planck 353 GHz dust polarization fraction. In this subsection, we discuss the assumptions made in previous studies regarding the Local Bubble surface. These assumptions help us design a test for studying this effect in the next subsection. A significant detection of an imprint of the Local Bubble geometry on the dust polarization fraction would validate these assumptions.
@@
The first assumption is that the observed polarized dust emission is dominated by dust in the Local Bubble wall at the relevant angular scales and Galactic latitudes. This assumption, made in several analyses (e.g., Alves et al., 2018; Pelgrims et al., 2020; O’Neill et al., 2023), is supported by several studies using optical starlight polarization data (Leroy, 1999; Andersson & Potter, 2006; Santos et al., 2011; Frisch et al., 2015; Medan & Andersson, 2019; Cotton et al., 2019; Skalidis & Pelgrims, 2019). The alignment of neutral hydrogen structures at local velocities with starlight polarization toward stars at distances within a few hundred parsecs is consistent with a picture where most of these structures are positioned at comparable distances within the Local Bubble at high Galactic latitudes ($|b|~{}>~{}30^{\circ}$) (Clark et al., 2014). Gontcharov & Mosenkov (2019) observed that starlight polarization fraction plateaus after 150-250 pc across the sky. By comparing the 353 GHz polarized emission with the polarized optical starlight, Skalidis & Pelgrims (2019) find that most of the 353 GHz polarized emission signal is captured within the first 250 pc at $|b|~{}>~{}60^{\circ}$, suggesting the presence of a dust wall around that distance. They, however, find that this conclusion does not hold at $30^{\circ}~{}<~{}|b|~{}<~{}60^{\circ}$.
@@
The second assumption is that the magnetic field’s inclination is tangential to the surface of the Local Bubble, which stems from the model assumed for the formation of the Local Bubble. To fit a model of the Local Bubble magnetic field, Alves et al. (2018) and Pelgrims et al. (2020) assume that all the swept-up matter and field lines due to the supernova explosions that formed the Local Bubble are squeezed into a thin layer that follows its surface, leading the magnetic field lines to be tangent to the surface. O’Neill et al. (2023) also make this assumption to project the observed polarization angles of the dust emission onto the Bubble’s surface and build a 3D model of the Bubble wall magnetic field.
@@
Other works have found magnetic field structure tangential to bubbles on supernova scales, observationally (e.g., Kothes & Brown, 2009; West et al., 2016; Tahani et al., 2022a, b) and in simulations (Kim & Ostriker, 2015; Maconi et al., 2023). It has also been shown on the scales of Hii regions, observationally (Tahani et al., 2023) and in simulations (Krumholz et al., 2007). On the scale of superbubbles, such as our Local Bubble, a comparison of the plane-of-the-sky and line-of-sight magnetic field strengths as well as measurements of the dust polarization fraction towards regions associated with the Orion-Eridanus superbubble suggest that the large-scale magnetic field in the region was primarily shaped by the expanding superbubble and is tangential to its surface (Heiles, 1997; Soler et al., 2018). However, there is no direct evidence that the nearby magnetic field is preferentially tangential to the Local Bubble surface.
@@
The shape of the Local Bubble wall has been modeled differently in different works. Some studies fit a generalized parametric geometry, such as an ellipsoid (Alves et al., 2018), in an attempt to fit general properties of the Local Bubble magnetic field on large scales. Others model the detailed boundary of the Local Bubble using 3D maps of the dust extinction (Pelgrims et al., 2020; O’Neill et al., 2024). These models therefore vary based on both the variations in the different 3D dust maps used as well as the methodology applied to these maps to define a Local Bubble surface. Pelgrims et al. (2020) model the radial distance of the Local Bubble wall from the Sun in each direction as the first distance where the second derivative with respect to the distance of the differential extinction constructed by Lallement et al. (2019) reaches zero, i.e., the first inflection point, ${\rm d}^{2}A_{V}^{\prime}(r)/{\rm d}r^{2}~{}=~{}0$. O’Neill et al. (2023) and Zucker et al. (2022) use the geometry defined by Pelgrims et al. (2020) in their work. O’Neill et al. (2024) employ a similar methodology to Pelgrims et al. (2020), using the differential extinction maps constructed by Edenhofer et al. (2023) instead to construct their model. Other studies use different tracers to model the geometry of the Bubble. Liu et al. (2017) assumes that the measured X-ray intensity is proportional to the distance to the Bubble in the considered direction. Several other tracers such as NaI absorption measurements (Sfeir et al., 1999; Lallement et al., 2003), stellar color excess measurements (Lallement et al., 2014), and diffuse interstellar bands (Farhang et al., 2019), have also been used for constructing models of the Bubble wall geometry. These geometries vary significantly from one to another. While some model the Local Bubble as a closed surface (Pelgrims et al., 2020), others describe the same structure as a Local Chimney, i.e., open in one or both directions away from the disk and funneling material into the Milky Way’s halo (e.g., Heiles, 1984; Sfeir et al., 1999; Lallement et al., 2003; Marchal & Martin, 2023; O’Neill et al., 2024). Also, the Local Bubble surface may have tunnels to surrounding cavities like the Gum Nebula and/or GSH238+00+09 rather than having a closed geometry (e.g., Welsh, 1991; Lallement et al., 2003; Marchal & Martin, 2023; O’Neill et al., 2024).